Print

Print


Ticket Number: HEA-NOC/20080221-1      Ticket Status: UPDATE
Ticket Type:   unscheduled             Resolver:      
Ticket Opened: 20080221 09:44 UTC      Problem Start: 20080220 21:05 UTC
Ticket Update: 20080225 12:36 UTC      Problem End:   

Site/line:     St. Angela's

Problem Description:
Circuit to St. Angela's flapping repeatedly from 21:05 last night until
7:35am.

Problem Effects:
Periodic and repeated loss of connectivity to St. Angela's.

Problem History:
20080221 09:17 GAR Cleared interface counters (recording existing values in
extra ticket info field)
After 10 minutes, interface shows 21 input errors and 24 CRC errors.
Raising to BT.

20080221 09:52 GAR Call received overnight from BT, regarding this link:
Call Date/Time: 21/02/2008 02:41:24
Name of institution:  BT NOC
Details of call:  WANTS NOC TO CHECK HEAN0122
BT Ticket Ref: TICKET 1-562154065

Confirmed with BT that endpoint routers were up and running, and they will
continue their investigation.

Ticket held open waiting for feedback from BT engineers.

20080222 09:58 DW Further flaps this morning; called BT helpdesk, who are
awaiting an update from the engineer assigned to the case, and expect
updates every two hours. This call has been logged by the as a P2 - have
asked them to record that the site is indeed down.

20080222 12:29 DW Called BT again. It's now upgraded to a P1 but still no
update - engineer will be asked to call us.

20080222 16:44 DW Called back. Engineer in charge is consulting with SDH
engineer regarding the radio part of the line. Still no estimate of time
to fix.

20080222 18:15 BN BT Ireland engineers will be investigating the link in
Sligo on Saturday 23rd.  No site access is available in St Angela's until
Monday morning at the earliest, so if site access is required they will
attend at that time, once this has been arranged with site contacts.  BT
Ireland will inform HEAnet when they have a more detailed ETA for the work
in Sligo.

20080224 12:17 BB BT have been checking the circuit on Saturday 23rd and
Sunday 24th.

20080225 09:36 DL call BT for update and their engineering department has
replicated the circuit in a lab environment to investigate the problem. No
estimated time to fix provided, will provide update along this morning.

20080225 10:23 DL Call BT NMC, replication circuit in lab is running
stable and at the moment is to operate this way for a period of time
(unknown by NMC) until engineering determines the systems are stable. Once
stability is determined, the system is to be sent to Sligo to replace the
existing equipment. It is BT intention to replace theses systems along
today since the weather forecast announces weather conditions not suitable
for wireless circuits. In any case NMC cannot ascertain that this will
happen.

20080225 12:35 DL BT NMC informs test kit is running clean now and being
dismantled to go to Sligo this afternoon

Time to Fix:
Not yet known.

-- 
This ticket can be monitored at http://www.hea.net/tickets/20080221-1

HEAnet Limited. Registered in Ireland, No. 275301.